# Airdrops: Giving Money Away Is Harder Than It Seems

### v1

@@
Abstract.
Airdrops are used by blockchain applications and platforms to attract an initial user base, and to grow the user base over time. In the case of many airdrops, tokens are distributed to select users as a “reward” for interacting with the underlying platform, with a long-term goal of creating a loyal community that will generate genuine economic activity well after the airdrop has been completed. Although airdrops are widely used by the blockchain industry, a proper understanding of the factors contributing to an airdrop’s success is generally lacking.
@@ Abstract.
1. Introduction
@@
Blockchain platforms commonly create reward schemes to lure new customers and maintain the loyalty of existing users. In recent years, giving platform-minted tokens to users, or airdropping tokens, has proven itself as a particularly popular scheme (Makridis et al., 2023). Although the use of airdrops is prevalent in the blockchain world, we present preliminary findings showing there is no significant correlation between performing an airdrop and a platform’s popularity relative to available alternatives. Intuitively, this is not optimal, and may result in wasting funds that can be used to improve the Quality-of-Service (QoS)offered by the platform.
While the basic premise of airdrops is relatively straightforward, the design space for such reward schemes is large, and may depend on the characteristics of the platform designing the airdrop. For example, some airdrop mechanisms target power-users and endow them with a generous amount of rewards in the hopes that these users will generate valuable economic activity, and thus attract other users as well.
@@
To understand why previous airdrops did not always achieve their goals and to quantify their success, we first articulate a set of reasonable desired outcomes for airdrops. Armed with a basic desiderata, we then go over previous airdrops, see how they measure up, and uncover some interesting insights in the process. Specifically, we analyze the data pertaining to four recent airdrops and conclude that the majority of proceeds (as high as 95%, typically only within a couple of transfer steps) have been sold via exchanges, implying that the airdrops fall short of their intended purpose and benefit mostly airdrop farmers as opposed to regular users. This comparison is followed by a description of common pitfalls faced by airdrops in the past. Given that the blockchain ecosystem is still quite young, it is to be expected that the theory and practice of blockchain airdrops is still in its infancy, and thus that previous airdrops may have not been entirely successful. Finally, we use our analysis to suggest potential ways to create better airdrop mechanisms which are more fair to honest users.
1.1. Contributions
Arbitrum study. We perform a comprehensive case study of the Arbitrum (Kalodner et al., 2018) airdrop by measuring elements such as transaction volume, token, distribution structure, and token value before and after the airdrop period.
Quantitative analysis. We quantitatively analyze airdrops executed by ENS (ENS, 2023), dYdX (Juliano, 2018), Gemstone (Haig, 2023; lingland09, 2023), and 1inch (Bukov and Melnik, 2020). We show that most of the funds received through these airdrops are sold on exchanges, instead of being used to create dapp or platform engagement on the part of end-users.
Qualitative analysis. We qualitatively analyze previous airdrops, and propose guidelines for developing future airdrops in an effort to avoid some of the pitfalls in the airdrops we analyze.
Scientific reproducibility. To ensure the scientific reproducibility of our empirical analysis, we share our dataset and scripts (Messias et al., 2023b). Refer Section A in the appendix for more details.
2. Airdrop Desiderata
Airdrops are typically used to bootstrap a community of users. We break down this overarching high-level goal into multiple sub-goals. These sub-goals are not necessarily orthogonal, and furthermore, there may be others that are worthwhile to pursue; we focus on them as they serve to illustrate interesting issues faced by common airdrop mechanisms.
1. Attract users in the short-term. Historically, airdrops were used by new blockchain platforms to create an initial user base. In particular, decentralized platforms commonly become more useful and attractive to users as a function of the amount of economic activity taking place.
2. Foster long-term involvement. Bootstrapping an initial base of users is important, but not enough for sustaining a high degree of economic activity over time. Ideally, platform users should become regular, everyday users. This can be achieved by issuing rewards that can only be used within the platform, similarly to frequent-flyer points. For example, in layer 2 blockchains (L2s) (Sguanci et al., 2021), one may award discounts on future transactions. Other measures that may help are having multiple airdrops, and rewarding users who perform “quests” that expose them to all that the platform has to offer. For example, Linea’s airdrop quests gave users an in-depth tour of the platform’s various features and use-cases (Labs, 2023a).
3. Target value-creating users. An airdrop should ideally focus on users who can contribute the most to the long-term sustainability of the platform. In the context of platforms that rely on user-based liquidity provisioning, this may refer to users who provide the most liquidity to lending pools (Yaish et al., 2022) and decentralized exchanges (DEXs) (Yaish et al., 2023), or across a variety of tokens. In rollups, one may consider so-called “creators” who deploy popular and useful contracts as particularly valuable. Such users provide additional use-cases for a platform, and thus serve to attract other users.
Table 1. Airdrops we analyze in this paper.
Airdrop Start date End date Blockchain
ENS October 2021 May 2022 Ethereum
dYdX September 2021 Still active Ethereum
1inch December 2020 Still active Ethereum
@@ 2. Airdrop Desiderata
3. Exchange Aftermarket Analysis
@@
The motivation for our quantitative analysis in this paper comes from the following observation: airdrop recipients quickly sell tokens and jump ship, which obviously defeats the point of airdrops in the first place. Analyses performed on DEXairdrops show that airdrop recipients sometimes sell all of their tokens shortly after receiving them. Thus, 75% of the recipients of Uniswap’s airdrop dumped all of their tokens within a week of receiving them (jhackworth, 2022), and 61% have sold their tokens shortly after the drop in the case of ParaSwap (Fan et al., 2023). In both cases, recipients mostly stopped using the corresponding platforms within months. This behavior indicates that the airdrop did not succeed in retaining recipients, or that recipients are mostly sybil accounts. Furthermore, a quick sell-off may alarm the market, if it is interpreted as a sign of a lack of belief in the long-term success of the platform.
Below we consider data related to the four airdrops (ENS, dYdX, Gemstone, and 1inch) in Table 1 that were collected from an Ethereum and zkSync Era archive nodes. We have made available both our data collection and analysis scripts, along with the dataset itself (Messias et al., 2023b). To identify exchanges, we use a list of 484 exchange account addresses on Ethereum’s mainnet obtained from Dune using the query:
select * from addresses_ethereum.dex
Platform Token Tokens % sold at # of recipient Tokens received per address
supply distributed exchanges addresses Mean Std. Median Min Max
ENS $1\times 10^{8}$ $1.96\times 10^{7}$ 36.51% $102822$ $190.93$ $141.57$ $152.4$ $0.0001$ $1143.54$
dYdX $1\times 10^{9}$ $1.77\times 10^{8}$ 35.35% $42682$ $2325.85$ $91336.75$ $392.56$ 00 $2.47\times 10^{7}$
1inch $1.5\times 10^{9}$ $8.3\times 10^{7}$ 53.47% $43402$ $1917.86$ $50164.49$ $636.48$ 00 $9.75\times 10^{6}$
Gemstone $5.06\times 10^{16}$ $5.04\times 10^{16}$ 95.01% $20014$ $2.31\times 10^{12}$ $6.93\times 10^{11}$ $2.51\times 10^{12}$ 0 $2.64\times 10^{12}$
Table 2. Airdrop distribution statistics for the four airdrops.
3.1. Tokens Distributed Over Time
Table 2 proves some numbers on how airdropped tokens are spread across the user base. We observe that the number of recipients is quite high for every analyzed airdrops, which may indicates some effort on the part of airdrop farmers (column 3). Moreover, the data indicates that the tokens were often traded on exchanges, as evident from the fact that the first transfer made by the addresses after the airdrop was a sale in an exchange (column 10). Gemstone stands out from the rest, with 95% of its tokens ending up on exchanges.
@@ 3.1. Tokens Distributed Over Time
Furthermore, Gemstone’s airdrop stands out significantly in terms of the number of tokens distributed compared to other airdrops. This substantial distribution also results in a notably higher median number of airdrops per recipient when compared to the other airdrops. Remarkably, Gemstone allocated the lion’s share, specifically 99.53%, of their total token supply during their airdrop campaign. It is important to highlight that Gemstone was executed primarily as a sybil attack rather than a legitimate airdrop.
Platform Airdrop recipients # of hops
Total # sent to exchanges % sent to exchanges Mean Std. Median Min. Max.
ENS $102822$ $86157$ 83.79% 1.75 1.06 2 1 50
dYdX $42682$ $41320$ 96.81% 2.25 0.59 2 1 18
1inch $43402$ $41913$ 96.57% 1.85 0.64 2 1 7
Gemstone $20014$ $20000$ 99.93% 1 1 1 1 1
@@
Table 3. Airdrop recipients’ interaction with exchanges. Gemstone stands out as an exception with a median of 1 hop, while other protocols show a median of 2 hops for connecting airdrop recipients with exchanges.
@@
Figure 1. Time for airdrop tokens to be transferred to exchanges.
Focusing on Table 3, we computed the shortest path taken from each recipient address of an airdrop to any of the exchange address in our set. Our findings reveal that the transfer of tokens to exchanges typically takes only a few steps, suggesting that airdrop farmers, in general, are not making significant efforts to hide their activities. The outlier is Gemstone where in all cases tokens were sold within only one hop. Surprisingly, the majority of accounts managed to reach an exchange with a relatively low number of intermediating hops. This observation underscores the prominent role that exchanges play within the crypto ecosystem.
@@
Figure 2. Network structure for selling airdropped tokens in 4 airdrops.
Figure 1 shows the CDFs that demonstrate how quickly tokens are sold at changes. Per Figure 0(a), we see that within about a million blocks, the majority of tokens are gone. For Gemstone the number of blocks significantly higher than others because of a delay introduced by airdrop farmers. As the block generation is quite different for Ethereum (a new block is created every 15s on average) and zkSync, we normalized the blocks in terms of the number of days. Per Figure 0(b), within a single day, 64.67% of the 1inch tokens were sent to an exchange. In comparison, ENS had a transfer rate of 45.86%, while dYdX stood at 37.41%. This does not align with one of the primary objectives of a protocol when conducting an airdrop, which is to promote sustained user engagement, as discussed in Section 2.
3.2. Token Distribution Graphs
To better understand the structure of the transfers that happened after each address received the airdrop tokens, we analyzed the transfer network. Denoted as G(V, E), every node (V) represents an address, and an edge (E) is established when tokens are transferred from one address to another. Specifically, the ENS network consists of $184585$ nodes and $608462$ edges, the dYdX network comprises $112853$ nodes and $406027$ edges, the Gemstone network encompasses $20003$ nodes and $240113$ edges, and the 1inch network features $308329$ nodes and $1400913$ edges.
To make these graphs more visually readable, we limited the number of hops in the data to one hop from any address that received the protocol airdropped the first few hours after the airdrop and plot their largest component in Figure 2. We manually labeled the higher in-degree nodes using labels provided by Etherscan, a popular blockchain explorer. We show that for all protocols but Gemstone, Uniswap is the decentralized exchange that received the majority of transfers (measured in terms of in-degree) from the addresses receiving the airdrops, followed by SushiSwap.

### v2

@@
Abstract
Airdrops are used by blockchain applications and protocols to attract an initial user base, and to grow the user base over time. In the case of many airdrops, tokens are distributed to select users as a “reward” for interacting with the underlying protocol, with a long-term goal of creating a loyal community that will generate genuine economic activity well after the airdrop. Although airdrops are widely used by the blockchain industry, a proper understanding of the factors contributing to an airdrop’s success is generally lacking.
@@ Abstract
keywords:
Airdrop design, Airdrop farmer, Web3, Layer 2
1 Introduction
@@
Blockchain protocols often design reward programs to attract new customers and foster the loyalty of existing users. In recent years, the practice of distributing platform-minted tokens, known as “airdropping,” has gained widespread popularity [53]. For example, in 2023 alone, 4.56 billion US dollars worth of tokens were airdropped to users across various protocols [9]. Although the use of airdrops is prevalent in the blockchain world, we present preliminary findings showing that there is no significant correlation between performing an airdrop and a platform’s popularity relative to available alternatives. Intuitively, this is not optimal and may result in the loss of funds that can be used to improve the Quality-of-Service (QoS)offered by the platform.
Although the fundamental concept of airdrops is relatively straightforward, the design space for such reward schemes is expansive and may vary depending on the characteristics of the platform implementing the airdrop. For instance, certain airdrop mechanisms focus on power users, granting them a substantial amount of rewards with the expectation that these users will stimulate valuable economic activity, subsequently attracting additional users. However, a potential drawback to this approach arises when the distributed tokens empower users to propose changes to a protocol through decentralized governance. It allows users to suggest or vote on changes within a platform, typically employing a one-token-one-vote strategy, where a single user may possess more than one voting token. This approach introduces the risk of voting power concentration, where a minority of users hold the majority of the decision-making voting power [56, 22, 68].
@@
To understand why previous airdrops did not always achieve their goals and to quantify their success, we first articulate a set of reasonable desired outcomes for airdrops. Armed with a basic desiderata, we then go over previous airdrops, see how they measure up, and uncover some interesting insights in the process. Specifically, we analyze the data pertaining to five popular airdrops (ENS [19], dYdX [39], 1inch [10], Arbitrum [41], and Uniswap [1]) and a fake airdrop conducted by a sybil farmer (Gemstone [35, 50]). We conclude that most of the proceeds (as high as 95%) have been sold via exchanges shortly after the airdrop, implying that the airdrops fall short of their intended purpose and benefit mostly airdrop farmers, highly-skilled users who employ sophisticated tactics to increase the share of airdropped tokens that they receive. Furthermore, we describe the common challenges faced by airdrops in the past. Given that the blockchain ecosystem is still quite young, it is expected that the theory and practice of blockchain airdrops are still in its infancy and thus that previous airdrops may not have been entirely successful in the long run. Finally, we use our analysis to suggest potential ways to create better airdrop mechanisms that are more fair to honest users.
1.1 Research Questions
In this section, we address three critical research questions central to our study.
@@ 1.1 Research Questions
We evaluate the effectiveness of blockchain protocols in converting short-term user incentives, such as airdrops, into long-term user engagement. The primary goal of conducting an airdrop is to attract users in the short-term and encourage them to become long-term users. Our investigation aims to measure how well these protocols achieve this objective.
Contrary to the assumption that airdrops are the primary driver of long-term user base growth, our research reveals otherwise. Through an exploration of past airdrops conducted, we demonstrate that airdrops may not be the sole determinant of sustained user retention. For example, findings from a recent report by Messari [8] indicate that users join a platform when short-term incentives are available, such as airdrops. However, these protocols struggle to retain users in the long-term, as users often shift their attention to protocols offering better incentives as soon as the initial rewards expire. This behavior implies that protocols successfully onboard users during incentive periods, like airdrops, but witness a drop in user engagement once the incentives come to an end.
Additionally, protocols not conducting airdrops also manage to onboard users due to an expectation of potential future airdrops. Our findings align with this observation, as our analysis indicates that, even after Arbitrum conducted an airdrop, other protocols attracted user attention and increased their total value locked (TVL). This suggests that airdrops alone may not be sufficient to ensure sustained user retention of the platform performing the airdrop, and other factors play crucial roles in long-term platform success.
[…]
Since protocols face challenges in achieving sustained user retention through airdrops alone, an alternative strategy involves incentivizing developers and innovators to build new applications, such as decentralized exchanges (DEXs)and non-fungible token (NFT). This approach aims to foster a dynamic ecosystem that benefits all users on the platform. To enhance user experience and maintain user loyalty, a focus on developing more, potentially exclusive, decentralized applications (dApps) is crucial. Another interesting approach is to provide additional incentives, such as discounts on fees for future platform interactions. For example, in the context of a rollup, users who receive airdropped tokens could redeem them for discounted fees, encouraging long-term engagement. Also, implementing multi-round airdrops allows protocols to learn from previous rounds, enabling iterative improvements to address previously identified issues [2]. Community engagement also plays a vital role in understanding the needs and preferences of users.
In line with these considerations, Blast has introduced an innovative airdrop strategy centered around user contributions to their protocol [55]. This strategy involves awarding points to users as a quantitative measure of their contribution. Points are granted for actions such as successfully bringing in new participants to the blockchain protocol, similar to a referral system, or for users who bridge (send) their tokens from one chain to Blast.
Furthermore, user engagement in blockchain protocols often hinges on short-term incentives. Users tend to join the platform when these incentives, such as airdrops, are available. However, a notable trend is the propensity of users to migrate opportunistically from one blockchain to another, following the trail of incentives. Consequently, a platform experiences an influx of users during incentive periods, only to witness a decline when these incentives cease.
[…]
A crucial aspect to consider involves airdrop farmers. These are sophisticated users who employ intricate tactics to maximize their chances of receiving more tokens per airdrop. They often create numerous wallet addresses controlled by a single entity, a practice known as Sybil [37, 64]. Detecting Sybils can be challenging [21, 51, 36, 73, 59], especially when these accounts execute automatic and randomized actions (e.g., token transfers, bridging, providing liquidity, exchanging tokens, NFTs, etc.) on the blockchain without linking their wallets. Furthermore, the accessibility of airdrop farming software facilitates the automated execution of such actions [70, 29, 12, 15]. Given its intricacy, current methods for identifying Sybils predominantly rely on graph network analysis [21, 51] and machine learning techniques [36, 73, 59].
Another common challenge involves airdropping governance tokens. These tokens empower holders to propose or vote on protocol changes. Distributing these tokens can be risky when users do not prioritize the protocol’s best interests. Additionally, it poses the danger of granting excessive power to certain users, as these protocols typically use a one-token-one-vote mechanism [56, 22, 68]. While acknowledging the challenges highlighted earlier, we believe that these challenges can serve as valuable lessons for protocols venturing into the realm of airdrops, offering insights into the potential path to success.
Furthermore, it is essential to uphold a positive reputation within the community by avoiding any occurrences of insider trading [5, 42]. Insider trading refers to individuals within an organization using privileged information to gain a financial edge over other protocol users. This practice is a common violation of securities laws in traditional financial markets [6], and engaging in it can lead to negative reactions from the blockchain community [69, 61, 23].
Finally, the points discussed above raise questions about whether protocols are effectively achieving their goals post their airdrop campaigns.
1.2 Summary of our Contributions
We summarize our contributions as follows.
[…]
2 Airdrop Desiderata
Airdrops are powerful tools for promoting protocols and acquiring users, attracting newcomers, and incentivizing existing users to engage with the blockchain protocol or application. They are widely utilized for these purposes, with numerous examples documented in the literature (see Table 1). Any protocol capable of creating tokens can utilize airdrops to distribute them to users. For example, blockchain rollup solutions like Arbitrum and Optimism, as well as DeFi applications such as Uniswap, 1inch, dYdX, and ENS have all employed airdrops.
Airdrops can manifest in various forms, with single-round and multi-round distributions being the most prevalent [2]. In a single-round airdrop, tokens are distributed to users all at once, whereas multi-round airdrops involve distributing tokens over multiple rounds each following different strategies. This approach allows leveraging insights from previous rounds to address challenges encountered, such as mitigating potential sybil (i.e., multiple accounts controlled by a single entity) by observing past user behavior patterns [2]. The choice between single-round and multi-round airdrops depends on the protocol’s objectives and community dynamics. Community engagement is a critical aspect of airdrop development; it should be carefully considered by protocol developers.
Moreover, the timing of an airdrop significantly impacts the number of eligible users receiving tokens. Established protocols opting for delayed airdrops may have a larger user base compared to newer projects executing airdrops sooner. This difference in user base size introduces an added layer of complexity when performing sybil detection, as there are more accounts to evaluate and potentially safeguard against. If not performed correctly, this could impact community happiness. In such instances, the protocol may inadvertently reward accounts linked to industrial farming, which could be viewed negatively by the community. However, addressing industrial farming poses significant challenges and complexities. To mitigate this issue, LayerZero Labs implemented a self-reporting mechamism for sybils. Under this system, sybils have the option to self-report and receive 15% of their designated token allocation. Failure to self-report may result in them not receiving tokens if they are identified by the company or its affiliates [46].
Table 1: List of popular airdrops we analyze in this paper.
[…]
dYdX Sept. 2021 Still active Ethereum Continuous DEX
1inch Dec. 2020 Still active Ethereum Continuous DEX
[…]
Arbitrum Mar. 2023 Sept. 2023 Arbitrum One Single Single
Uniswap Sept. 2020 Still active Ethereum Continuous DEX
[…]
Next, we break down airdrop’s high-level goal to bootstrap a community of users into multiple sub-goals. These sub-goals are not necessarily orthogonal, and furthermore, there may be others that are worthwhile to pursue; we focus on them as they serve to illustrate interesting issues faced by common airdrop mechanisms.
Attract users in the short-term.  Historically, airdrops have been employed by new blockchain protocols to establish an initial user base and to provide an initial boost of liquidity to the underlying chain and its protocols. Decentralized platforms, in particular, tend to become more useful and attractive to users as the level of economic activity increases.
Foster long-term involvement.  Bootstrapping an initial base of users is important, but not enough for sustaining a high degree of economic activity over time. Ideally, users should become regular, everyday users. This can be achieved by issuing rewards that can only be used within the blockchain protocol or application, similarly to frequent-flyer points. For example, in layer-2 blockchains (L2s) [67], one may award discounts on future transactions. Other measures that may help are having multiple airdrops, and rewarding users who perform “quests” that expose them to all that the blockchain has to offer. For example, Linea’s airdrop quests gave users an in-depth tour of their various features and use-cases [44].
Target value-creating users.  An airdrop should ideally focus on users who can contribute the most to the long-term sustainability of their platform. In the context of platforms that rely on user-based liquidity provisioning, this may refer to users who provide the most liquidity to lending pools [78] and DEXs [77], or across a variety of tokens. In rollups, one may consider so-called “creators” who deploy popular and useful contracts as particularly valuable. Such users provide additional use cases for a platform, and thus serve to attract other users.
3 Exchange Aftermarket Analysis
@@
The motivation for our quantitative analysis in this paper comes from the following observation: airdrop recipients quickly sell tokens and jump ship, which obviously defeats the point of airdrops in the first place. Analyses performed on DEXairdrops show that airdrop recipients sometimes sell all of their tokens shortly after receiving them. For example, 75% of the recipients of Uniswap’s airdrop dumped all of their tokens within a week of receiving them [38], and 61% have sold their tokens shortly after the drop in the case of ParaSwap [21].
In both instances, the majority of recipients ceased using the relevant blockchain protocols within a few months. This pattern suggests that the airdrop was ineffective in keeping recipients engaged, or that many recipients were likely sybil accounts. Additionally, a rapid sell-off could unsettle the market, particularly if seen as an indication of dwindling confidence in the protocol’s future prospects. Here, we examine data associated with six airdrops (ENS, dYdX, 1inch, Arbitrum, and Uniswap, including Gemstone) presented in Table 1, which were gathered from archive nodes of Ethereum and zkSync Era.To identify exchanges, we use a list of 484 exchange account addresses on Ethereum’s mainnet obtained from Dune using the query: select * from addresses_ethereum.dex and from Etherscan [20].
Table 2: General airdrop distribution statistics.
[…]
supply distributed exchanges addresses
[…]
dYdX $1\times 10^{9}$ $1.77\times 10^{8}$ 35.45% $42682$
1inch $1.5\times 10^{9}$ $83.2\times 10^{6}$ 54.05% $43402$
Gemstone $50.6\times 10^{15}$ $50.4\times 10^{15}$ 95.01% $20014$
[…]
Table 3: Tokens received per address.
[…]
dYdX $2325.85$ $91336.75$ $392.56$ 00 $24.7\times 10^{6}$
1inch $1917.86$ $50164.49$ $636.48$ 00 $9.75\times 10^{6}$
Gemstone $2.31\times 10^{12}$ $6.93\times 10^{11}$ $2.51\times 10^{12}$ 0 $2.64\times 10^{12}$
Arbitrum $1874.02$ $1566.92$ $1250$ $625$ $10250$
Uniswap $620.13$ $9347.91$ $400$ $400$ $2.1\times 10^{6}$
[…]
3.1 Tokens Distributed Over Time
Table 2 presents the number of airdrops distributed and subsequently sold on exchanges. Additionally, Table 3 details the distribution of airdrops per user. These tables provide quantitative insights into how airdropped tokens are disseminated across the user base. Our analysis reveals a significant number of recipients for each airdrop, suggesting potential involvement of airdrop farmers (see column 5 in Table 2). Furthermore, the data indicate that tokens were frequently traded on exchanges. This is supported by the observation that the initial transfer made by the addresses post-airdrop was often a sale on an exchange (see column 4 in Table 2). Notably, the case of Gemstone is particularly striking, with 95% of its tokens being sold on exchanges. In this particular Gemstone case, a non-open sourced DEX created by the airdrop farmer [35].
@@ 3.1 Tokens Distributed Over Time
Furthermore, Gemstone’s airdrop stands out significantly in terms of the number of tokens distributed compared to other airdrops. This substantial distribution also results in a notably higher median number of airdrops per recipient when compared to the other airdrops as shown in Table 3. Remarkably, Gemstone allocated the lion’s share, specifically 99.53%, of their total token supply during their airdrop campaign. It is important to highlight that Gemstone was executed primarily as a sybil attack rather than a legitimate airdrop.
Figure 1: Tokens claimed per day for each of the Airdrops analyzed.
Among all the airdrops analyzed in this study, the Arbitrum airdrop exhibited the fastest rate of token claims by users. Figure 1 illustrates the distribution of accounts claiming their tokens each day. Arbitrum conducted a substantial airdrop, distributing $1162166000$ ARB tokens among $625143$ selected accounts. Notably, $583137$ (93.28%) of these accounts successfully claimed 94.03% of the allocated ARB tokens. Remarkably, 72.45% of the accounts completed their token claims on the first day, with an additional 14.41% doing so on the following day. This cumulative figure amounts to almost 87% of accounts claiming the Arbitrum airdrop within one day of its release, indicating high user engagement and prompt action among the majority of participants.
Table 4: Airdrop recipients’ interaction with exchanges.
[…]
Total # sent to exchanges % sent to exchanges
[…]
dYdX $42682$ $41320$ 96.81%
1inch $43402$ $41913$ 96.57%
Gemstone $20014$ $20000$ 99.93%
Arbitrum $583137$ $555040$ 95.18%
[…]
@@
Table 5: Distribution of hops for airdrop recipients connecting with exchanges.
[…]
Mean Std. Median Min. Max.
[…]
dYdX 2.25 0.59 2 1 18
1inch 1.85 0.64 2 1 7
Gemstone 1 1 1 1 1
Arbitrum 2.71 $0.79$ $3$ $1$ $32$
Uniswap 2.11 $0.43$ 2 1 9
[…]
@@
Figure 2: Time for airdrop tokens to be transferred to exchanges.
Users typically interact with exchanges to convert one token for another and to sell them. To assess how often users receiving an airdrop sell these tokens on exchanges for profit, we analyzed user interactions with exchanges post-airdrop. Table 4 indicates that a significant proportion of airdrop recipients engage with exchanges to trade their tokens. This engagement ranges from a minimum of 83.79% in the case of ENS to a maximum of 99.93% for Gemstone.
Additionally, Table 5 presents the shortest path from each recipient address of an airdrop to any exchange address in our dataset. Our findings reveal that the transfer of tokens to exchanges generally involves only a few steps, suggesting that airdrop recipients do not make significant efforts to conceal their activities. Gemstone is a notable outlier, with all tokens being sold within a single hop. Surprisingly, most accounts managed to reach an exchange with a relatively low number of intermediary hops (between 2 and 3 hops). This observation underscores the pivotal role that exchanges play within the cryptocurrency ecosystem.
@@
(e) Arbitrum.
(f) Uniswap.
Figure 3: Network structure for selling airdropped tokens in 6 airdrops.
Figure 2 shows the cumulative distribution functions (CDFs) demonstrating how quickly users interact with an exchange right after receiving tokens from an airdrop. As shown in Figure 2(a), the majority of accounts exchanged their tokens within approximately one million blocks. The number of blocks for Gemstone is significantly higher than others due to a delay introduced by airdrop farmers. Given the difference in block generation times between Ethereum (with a new block created approximately every 15 seconds) and zkSync, we normalized the blocks in terms of days. According to Figure 2(b), within a single day, 64.67% of the accounts interacted with an exchange. In comparison, ENS had an interaction rate of 45.86%, while dYdX stood at 37.41%. This rapid exchange behavior does not align with one of the primary objectives of a protocol conducting an airdrop, which is to promote sustained user engagement, as discussed in Section 2. The prompt exchange of tokens suggests that users might be leaving a protocol soon after receiving their airdrops.
3.2 Token Distribution Graphs
To better understand the structure of the transfers that happened after each address received the airdrop tokens, we analyzed the transfer network. Denoted as G(V, E), every node (V) represents an address, and an edge (E) is established when tokens are transferred from one address to another. Specifically, the ENS network consists of $184585$ nodes and $608462$ edges, the dYdX network comprises $112853$ nodes and $406027$ edges, the Gemstone network encompasses $20003$ nodes and $240113$ edges, and the 1inch network features $308329$ nodes and $1400913$ edges. Lastly, Arbitrum accounts for $921650$ nodes and $27438608$ edges while Uniswap contains $1180830$ nodes and $3762613$ edges.
To make these graphs more visually readable, we limited the number of hops in the data to one hop from any address that received the protocol airdrop the first few hours after the airdrop and plot their largest component in Figure 3. We manually labeled the higher in-degree nodes using labels provided by Etherscan, a popular blockchain explorer [20]. We show that for all protocols but Gemstone, the decentralized exchange that received the majority of transfers (measured in terms of in-degree) from airdrop recipients is Uniswap, followed by SushiSwap.